IT is approximately 480 miles from Los Angeles, California to San Francisco, Californis. Allowing for various traffic conditions, a driver can average approximately 60 miles per hour.
a)I need to write a linear function that expresses the distance traveled, d, as a function of time,t.
b) How far have you traveled in 4 hours?
c)Write a linear function that expresses the distance to be traveled to reach San Francisco , s ,as a function of time, t.
d) How far will I need to travel to reach San Francisco after I have traveled 3 hours?
Ive been out of school for 21 years and never had to take algenbra and this is very hard for me to understand this question so I can help my daughter in her Algebra class. Please Help!!!!!!!!!

Answer by cowess(34)

a) d=60t
b) 4*60=240
c) s=480-60t
d) 480-60*3=480-180=300